# Continual Low-Rank Scaled Dot-product Attention

### v2

@@ I Introduction
Each of the three input matrices in Eq. 1 has a dimensionality of $n\times d$, where $n$ is the number of tokens and $d$ the number of features, leading the computational and memory costs of the SDA module to have a quadratic form $O(n^{2}d)$. For large $n$, this attention mechanism can be too expensive, both computationally and memory-wise, for certain tasks where model inference needs to be performed with limited resources, such as tasks coming from applications in Robotic Perception and Control [6, 7], Forecasting [8], and Computer Vision [9, 10].
@@
While efficient model architectures have been proposed that can lead to fast inference for static tasks, such as the classification of images, videos or audio sequences, Continual Inference tasks remain challenging for Deep Learning models (even for those models that can operate in real time on the corresponding static task). Continual Inference [27] can be defined as the process of providing a result for each input of a continual data stream, for instance performing event/action classification based on the visual frames captured by a camera operating continually for a long period of time. An inherent requirement for solutions targeting Continual Inference tasks is that they need to be able to process the incoming stream data with low latency and low resource consumption.
@@
The adoption of Transformer models, including the original and approximate formulations as well as the efficient architectures mentioned above, in a Continual Inference setting requires a transformation of the stream data processing task to its corresponding static task every time inference is performed. This is usually achieved with a sliding temporal window that creates a sequence segment (e.g., when processing visual streams, a video clip) formed by the newly captured data frame and the data frames preceding it. The full sequence within the window is then passed through the model to provide an inference result. This process leads to a very high amount of redundant computations when performing inference over successive windows, as every input data frame needs to be processed more than once (specifically, $n$ times for a sliding window with size $n$ data frames and step size of one). Continual Transformers [28] were proposed for addressing this limitation by processing each input token once. This is done by caching and reusing intermediate results to eliminate redundant computations in sequence data processing, following the principles of Continual Inference Networks (CINs) [27, 28, 29, 30]. The Continual Transformers achieve a linear computational and memory cost in the order of $O(nd)$.
@@
In this paper, we propose a new formulation of the Continual Transformer which further improves its memory and computational cost requirements for Continual Inference settings. We incorporate the Nyström-based approximation of the matrix multiplication followed by softmax nonlinearity in SDA, to further lower the Continual SDA’s memory footprint and the number of computations compared to its original formulation. To do this, we derive the model updates of the Nyström-based SDA in a continual manner. We propose two different ways to determine the landmarks used for processing continual stream data in the SDA approximation, and make the corresponding model architectural modifications.
@@ III-A Continual Transformer Attention
In the Continual Retroactive Attention, new tokens update the attention values of all the previous tokens within the attention window. All of these changes allow us to update the attention with $513nd+8n-9d-$ 5 operations. The peak memory cost (i.e., the maximum memory required during an inference step) of the update is $5nd+2n$ elements, while the valley memory cost (i.e., the memory cost between two successive inference steps) is $44nd+n-d-$ 4 elements. If only the newest token inference is needed, the Single Output Attention can be used to save computations and memory space, i.e.:
@@
where $a=\rho(q_{\text{new}},K)$. This formulation brings the number of operations down during inference to $14nd+n+d-$ 1, while the peak memory cost of the update is $2nd+n+2d$ elements and the valley memory cost is $22nd-$ 2 elements.
@@ III-B Nyström-based Attention
where $\Omega^{\dagger}$ is the Moore-Penrose pseudo-inverse of matrix $\Omega$, and $\tilde{Q}$ and $\tilde{K}$ are matrices formed by sets of $m$ landmarks, computed as the segment-means of the matrices $Q$ and $K$, respectively. This approach leads to an (approximate) attention mechanism with a computational cost of $4ndm+2nd+nm^{2}+2nm+dm^{2}+24m^{3}+23m^{2}$, peak memory cost of $14nd+2nm+2dm+6m^{2}+m+$ 1 elements and valley memory cost of $33nd-$ 3 elements. When $m\ll n$, it leads to a large reduction of both costs compared to the standard SDA formulation in Eq. (1), i.e., $O(nd)$ computational and memory costs. The same formulation has been used in [33] where, instead of defining the landmarks as the segment-means, landmarks are chosen from the sequence incrementally, in such a way that at each step the most orthogonal vector to the already selected tokens is chosen. In [34], attention is computed over the parts of the image of a video where the most elements have changed, triggering $m$ updates over the image as the rest of the weights are re-used.
@@ IV Method
As can be seen, in Eqs. (9)-(12), the landmark matrices $\tilde{Q}$ and $\tilde{K}$ are involved in the calculation of in all three matrices $B$, $\Gamma$ and $\Delta$. The Nyströmformer [19] calculates new landmarks for every inference step. However, this approach would lead to computational redundancies in a Continual Inference setting, as landmarks would need to be fully recomputed after every inference step. To address this issue, we exploit properties stemming from the fact that successive inference steps involve processing of highly-overlapping sequence data. We propose two ways for landmark selection, leading to model updates described in Sections IV-A and IV-B. Considerations related to the implementation aspects of the proposed model and its updates are discussed in Section IV-C.
Figure 1: Continual landmarks calculation (for $n=20$ and $m=4$). Until enough new input tokens have been received, the landmarks remain fixed. This includes the landmark calculated by some old tokens which have been shifted out of the current attention window (top). When enough input tokens have been received (bottom), a new landmark is computed using all the new tokens, replacing the $\tilde{q}_{\text{old}}$ landmark. The same approach is applied to update $\tilde{K}$.
@@ IV-A Continual Landmarks
Following the process of receiving new input tokens updating the matrices $Q$ and $K$ in a continual manner, the landmark matrices $\tilde{Q}$ and $\tilde{K}$ can be updated periodically, i.e., after a pre-defined number of input tokens are received, as illustrated in figure 1. This means that the landmark matrices $\tilde{Q}$ and $\tilde{K}$ are updated after receiving $\frac{n}{m}$ input tokens, following the segment-means process used in [19]. In practice, this will cause the oldest landmarks $\tilde{q}_{\text{old}}$ and $\tilde{k}_{\text{old}}$ (computed by using tokens that have already been shifted out of the $Q$ and $K$ matrices) to be shifted out of the landmark matrices $\tilde{Q}$ and $\tilde{K}$, respectively. The newest landmarks $\tilde{q}_{\text{new}}$ and $\tilde{k}_{\text{new}}$, computed as the mean of the most recent $\frac{n}{m}$ tokens in the $Q$ and $K$ matrices, respectively, will then be included in the landmark matrices $\tilde{Q}$ and $\tilde{K}$. As such, the update of the landmark matrices takes the form:
@@ IV-A1 Continual Inference with updated landmarks
where $B_{\phi}=(\phi(B)^{-1}\odot B)$ and $\Gamma_{\phi}=(\phi(\Gamma)^{-1}\odot\Gamma)$. Considering the need to store the intermediate matrices to perform the updates when there is no change in the landmarks, the total computational cost of Continual Inference when landmarks are updated is $8nd+nm^{2}+11n+17dm+24m^{3}+22m^{2}+23m$.
Following the Continual Transformer approach [30], a Single Output version of the Continual SDA is also proposed, as there are cases where only the attended features from the newest input token are needed, allowing for further reducing the number of computations. In practice, this is achieved by using the last row of the matrix $B_{\phi}$ rather than the full matrix i.e.,:
@@
where $(\beta_{\phi})_{\text{new}}$ is the last row of the $B_{\phi}$ matrix. This still requires performing the common intermediate computations in the same way as its retroactive counterpart. The total computational cost of these is $ndm+3nd+n+14dm+24m^{3}+23m^{2}+20m$.
@@ IV-A2 Continual Inference with non-updated landmarks
where $\phi(\Delta)^{-1}_{\text{prev}}$ and $(\Delta V)_{\text{prev}}$ correspond to the matrices obtained in the previous inference step. The total computational cost of Continual Inference when landmarks are updated is $ndm+8dm+m^{2}+8m$.
@@
The total computational cost of the Single Output version is reduced to $8dm+m^{2}+6m$.
Regarding the memory cost, for the Retroactive Continual Nyströmformer with continual landmarks, the valley and peak memory costs are $63nd+2nm+4dm+2m^{2}-2m-$ 6 and $4nd+3nm+2n+4dm+7m^{2}+4m$, respectively. For the Single Output Continual Nyströmformer with continual landmarks, the valley and peak memory costs are $22nd+4dm+2m^{2}-$ 2 and $2nd+nm+n+4dm+d+7m^{2}+4m$, respectively.
@@ IV-B Fixed Landmarks
After training the model and determining all landmarks, the SDA module used for Continual Retroactive Inference, i.e., $\text{Att}_{\text{CoNyRe}}^{\text{Fix}}$, has the form of Eq. (23), and the SDA module for the Continual Single Output Inference, i.e., $\text{Att}_{\text{CoNySi}}^{\text{Fix}}$, has the form of Eq. (26). The computational cost of this model is identical to the Nyström-based Continual Inference with non-updated landmarks for both the Retroactive and Single Output versions. Regarding the memory costs, for the Retroactive Continual Nyströmformer with fixed landmarks, the valley memory cost is $nm+3dm+m^{2}+m$ and the peak cost is $nd+nm+3dm+m^{2}+2m$. For the Single Output Continual Nyströmformer with fixed landmarks, the valley and peak memory costs are $3dm+m^{2}+m$ and $3dm+d+m^{2}+2m$, respectively.
@@ IV-C Implementation Aspects
  • For training, we use a modified version of the non-continual model with the circular positional encoding described above and the corresponding landmark selection scheme as described in Sections IV-A and IV-B, depending on whether continual or fixed landmarks are used, respectively. We follow this approach as the non-continual training processes are faster, since the entire sequence is available from the beginning and both continual and non-continual SDA variants produce identical results.
@@
For replacing the softmax with exponential operations in both the Continual Transformers and our Continual Nyströmformers numerical stability issues need to be considered, as the calculation of the exponential can be prone to overflow and underflow. This is addressed in most softmax implementations (e.g., [38]) by employing the so-called stable softmax variant [39] $\text{s}(x)_{i}=\text{exp}(x_{i}-C)/\sum_{j=1}^{n}\text{exp}(x_{j}-C)$. By setting $C$ to the maximum value in $x$ possible overflow and underflow issues are addressed, as at least one value will be higher than zero after the calculation of the exponential operations. However, this approach cannot be applied in the continual versions of SDA, as the maximum values of $x$ in the attention window can change at every inference step. This would cause constant updates in most matrices, increasing the cost very significantly. In our experiments described in the next section, this issue has not been observed. In case such stability issue is observed, using a dropout layer or other type of normalization layer before the SDA module can address it.
@@ V Experiments
In this Section, we provide comparisons of models using the proposed SDA formulations with models using the original SDA formulation [1], the Nyström-based SDA formulation [19] and the Continual SDA formulation [28] in terms of their memory and computational costs, as well as their performance on Audio Classification and Online Action Detection tasks. A summary of the properties of these models is provided in Table I.
TABLE I: Properties of the SDA modules used in our experiments.
@@
$\text{Att}_{\text{CoNy}}^{\text{Cont}}$checkmark $\checkmark$ Continual
$\text{Att}_{\text{CoNy}}^{\text{Fix}}$checkmark $\checkmark$ Fixed
@@ V-B Memory overhead experiments
The rest of the SDA formulations exhibit a similar behavior, as all of them require storing intermediate matrices with size $n\times d$, including the full matrices $Q$, $K$ and $V$. The SDA formulations $\text{Att}_{\text{CoNy}}$ and $\text{Att}^{\text{Fix}}_{\text{CoNy}}$ have the highest memory cost in this group, because they require storing the same set of matrices at peak memory time. The SDA formulations $\text{Att}_{\text{Co}}$ and $\text{Att}^{\text{Cont}}_{\text{CoNy}}$ have a very similar memory costs. This is due to that both SDA formulations need to store the same number of matrices of size $n\times d$ at peak memory time.
TABLE II: Audio classification on the GTZAN dataset.
@@
Att 1 $93.12\pm 5.44$ 11.1 $\times$1
2 $94.32\pm 4.31$ 22.19 $\times$1
$\text{Att}_{\text{Co}}$ 1 $93.16\pm 5.39$ 0.09 $\times$123.3
2 $94.52\pm 4.05$ 0.39 $\times$56.90
$\text{Att}_{\text{Ny}}$-4 1 93.54±$\pm 4.26$ 0.42 $\times$26.43
2 93.92±$\pm 4.43$ 0.85 $\times$26.11
$\text{Att}_{\text{Ny}}^{\text{Fix}}$-4 1 $93.24\pm 5.30$ 0.42 $\times$26.43
2 $93.76\pm 5.14$ 0.84 $\times$26.42
$\text{Att}_{\text{CoNy}}^{\text{Cont}}$-4 1 $92.92\pm 5.43$ 0.01 $\times$1110
2 $93.54\pm 4.91$ 0.12 $\times$184.9
$\text{Att}_{\text{CoNy}}^{\text{Fix}}$-4 1 $90.76\pm 5.52$ 0.006 $\times$1850
2 $92.36\pm 5.52$ 0.10 $\times$221.9
TABLE III: Online Action Detection on the THUMOS14 dataset using a Temporal Segment Network pre-trained on the Kinetics-400 dataset or ActivityNet dataset.
@@
Att 1 $64.32\pm 0.66$ $98.27\pm 0.05$ $56.00\pm 0.71$ $97.39\pm 0.04$ 16.85 $\times$1
2 $\textbf{64.66}\pm 0.30$ $98.28\pm 0.06$ $\textbf{56.95}\pm 0.80$ $97.36\pm 0.06$ 33.69 $\times$1
$\text{Att}_{\text{Co}}$ 1 $63.89\pm 0.14$ $98.33\pm 0.05$ $55.69\pm 0.41$ $\textbf{97.43}\pm 0.04$ 0.26 $\times$64.81
2 $63.93\pm 0.43$ $\textbf{98.34}\pm 0.02$ $56.04\pm 0.27$ $97.40\pm 0.08$ 1.11 $\times$30.35
$\text{Att}_{\text{Ny}}$-16 1 $59.74\pm 0.29$ $97.72\pm 0.03$ $50.57\pm 0.52$ $96.63\pm 0.05$ 4.71 $\times$3.58
2 $59.32\pm 0.23$ $97.72\pm 0.04$ $50.24\pm 0.14$ $96.47\pm 0.03$ 9.42 $\times$3.58
$\text{Att}_{\text{Ny}}^{\text{Fix}}$-16 1 $56.77\pm 0.51$ $97.46\pm 0.11$ $48.80\pm 0.75$ $96.41\pm 0.12$ 4.60 $\times$3.66
2 $53.80\pm 2.19$ $97.28\pm 0.29$ $48.53\pm 0.61$ $96.19\pm 0.11$ 9.21 $\times$3.66
$\text{Att}_{\text{CoNy}}^{\text{Cont}}$-16 1 $60.03\pm 0.33$ $97.70\pm 0.03$ $50.60\pm 0.13$ $96.64\pm 0.03$ 0.49 $\times$34.39
2 $59.30\pm 0.39$ $97.64\pm 0.07$ $50.68\pm 0.38$ $96.64\pm 0.04$ 1.82 $\times$18.51
$\text{Att}_{\text{CoNy}}^{\text{Fix}}$-16 1 $55.48\pm 0.26$ $97.21\pm 0.07$ $47.92\pm 0.69$ $96.05\pm 0.05$ 0.13 $\times$129.61
2 $52.41\pm 0.65$ $96.82\pm 0.13$ $46.60\pm 0.36$ $95.76\pm 0.06$ 1.31 $\times$25.71

### v3

@@ I Introduction
Each of the three input matrices in Eq. (1) has a dimensionality of $n\times d$, where $n$ is the number of tokens and $d$ the number of features, leading the computational and memory costs of the SDA module to have a quadratic form $O(n^{2}d)$. For large $n$, this attention mechanism can be too expensive, both computationally and memory-wise, for certain tasks where model inference needs to be performed with limited resources, such as tasks coming from applications in Robotic Perception and Control [6, 7], Forecasting [8], and Computer Vision [9, 10].
@@
While efficient model architectures have been proposed that can lead to fast inference for static tasks, such as the classification of images, videos or audio sequences, Continual Inference tasks remain challenging for Deep Learning models (even for those models that can operate in real time on the corresponding static task). Continual Inference [27] can be defined as the process of providing a result for each input of a continual data stream, for instance performing event/action classification based on the visual frames captured by a camera operating continually for a long period of time. An inherent requirement for solutions targeting Continual Inference tasks is that they need to be able to process the incoming stream of data with low latency and low resource consumption.
@@
The adoption of Transformer models, including the original and approximate formulations as well as the efficient architectures mentioned above, in a Continual Inference setting requires a transformation of the stream data processing task to its corresponding static task every time inference is performed. This is usually achieved with a sliding temporal window that creates a sequence segment (e.g., when processing visual streams, a video clip) formed by the newly captured data frame and the data frames preceding it. The full sequence within the window is then passed through the model to provide an inference result. This process leads to a very high amount of redundant computations when performing inference over successive windows, as every input data frame needs to be processed more than once (specifically, $n$ times for a sliding window with size $n$ data frames and step size of one). Continual Transformers [28] were proposed for addressing this limitation by processing each input token once. This is done by caching and reusing intermediate results to eliminate redundant computations in sequence data processing, following the principles of Continual Inference Networks (CINs) [27, 28, 29, 30].
Continual Inference methods are essentially redundant-free versions of their corresponding non-continual counterparts that produce identical outputs [28]. At the same time, the Nyström-based SDA formulation [19] has been extensively used in previous works [31] as an alternative to the Transformer with lower computational cost. The adaptation of this SDA formulation to a Continual Inference setting requires a comprehensive mathematical analysis. This can be used as a baseline for the continual adaptation of other low-rank approximation schemes, as the challenges are expected to be similar. In this paper, we make the following contributions:
@@
  • We propose a new formulation of the Continual Transformer, which further improves its memory and computational cost requirements for Continual Inference settings. We incorporate the Nyström-based approximation of the matrix multiplication followed by the softmax nonlinearity in SDA, to further lower the Continual SDA’s memory footprint and the number of computations compared to its original formulation. To do this, we derive the model updates of the Nyström-based SDA in a continual manner.
  • We propose two different ways to determine the landmarks used for processing continual stream data in the SDA approximation, and make the corresponding module modifications.
@@ III-A Continual Transformer Attention
In the Continual Retroactive Attention, new tokens update the attention values of all the previous tokens within the attention window. This allows the Continual Transformers to achieve linear computational and memory cost of $O(nd)$. If only the newest token inference is needed, the Single Output Attention can be used to save some additional computations and memory space, i.e.:
@@
where $a=\rho(q_{\text{new}},K)$.
@@ III-B Nyström-based Attention
where $\Omega^{\dagger}$ is the Moore-Penrose pseudo-inverse of matrix $\Omega$, and $\tilde{Q}$ and $\tilde{K}$ are matrices formed by sets of $m$ landmarks, computed as the segment-means of the matrices $Q$ and $K$, respectively. When $m\ll n$, it leads to a large reduction of both costs compared to the standard SDA formulation in Eq. (1), i.e., $O(nd)$ computational and memory costs. The same formulation has been used in [34] where, instead of defining the landmarks as the segment-means, landmarks are chosen from the sequence incrementally, in such a way that at each step the most orthogonal vector to the already selected tokens is chosen. In [35], attention is computed over the parts of the image of a video where the most elements have changed, triggering $m$ updates over the image as the rest of the weights are re-used.
@@ IV Method
As can be seen, in Eqs. (9)-(12), the landmark matrices $\tilde{Q}$ and $\tilde{K}$ are involved in the calculation of in all three matrices $B$, $\Gamma$ and $\Delta$. The Nyströmformer [19] calculates new landmarks for every inference step. However, this approach would lead to computational redundancies in a Continual Inference setting, as landmarks would need to be fully recomputed after every inference step. To address this issue, we exploit properties stemming from the fact that successive inference steps involve processing of highly-overlapping sequence data. We propose two ways for landmark selection, leading to model updates described in Sections IV-A and IV-B. Considerations related to the implementation aspects of the proposed model and its updates are discussed in Section IV-C. Table I shows the asymptotic computational and memory costs of the proposed method compared to the existing ones. A detailed analysis of this can be found in -A and -B.
TABLE I: Computational and memory cost of the different attention types. Valley and peak memory costs make reference to the lowest and highest memory costs of each attention type. A detailed analysis can be found in Apendices -A and -B.
SDA type Computational cost Valley memory cost Peak memory cost
Att $O(n^{2}d)$ $O(nd)$ $O(n^{2}+nd)$
$\text{Att}_{\text{Ny}}$ $O(ndm+nm^{2}+m^{3})$ $O(nd)$ $O(nd+nm+m^{2})$
$\text{Att}_{\text{Ny}}^{\text{Fix}}$ $O(ndm+nm^{2})$ $O(nd+m^{2})$ $O(nd+nm+m^{2})$
$\text{Att}_{\text{CoSi}}$ $O(nd)$ $O(nd)$ $O(nd)$
$\text{Att}_{\text{CoRe}}$ $O(nd)$ $O(nd)$ $O(nd)$
$\text{Att}_{\text{CoNySi}}^{\text{Cont}}$ (Ours) $O(ndm+m^{3})$ $O(nd+m^{2})$ $O(nd+m^{2})$
$\text{Att}_{\text{CoNyRe}}^{\text{Cont}}$ (Ours) $O(ndm+nm^{2}+m^{3})$ $O(nd+m^{2})$ $O(nd+nm+m^{2})$
$\text{Att}_{\text{CoNySi}}^{\text{Fix}}$ (Ours) $O(dm+m^{2})$ $O(dm+m^{2})$ $O(dm+m^{2})$
$\text{Att}_{\text{CoNyRe}}^{\text{Fix}}$ (Ours) $O(ndm+m^{2})$ $O(nm+m^{2})$ $O(nd+nm+m^{2})$
Figure 1: Continual landmarks’ calculation (for $n=20$ and $m=4$). Until enough new input tokens have been received, the landmarks remain fixed. This includes the landmark calculated by some old tokens which have been shifted out of the current attention window (top). When enough input tokens have been received (bottom), a new landmark is computed using all the new tokens, replacing the $\tilde{q}_{\text{old}}$ landmark. The same approach is applied to update $\tilde{K}$.
@@ IV-A Continual Landmarks
Following the process of receiving new input tokens updating the matrices $Q$ and $K$ in a continual manner, the landmark matrices $\tilde{Q}$ and $\tilde{K}$ can be updated periodically, i.e., after a pre-defined number of input tokens are received, as illustrated in Figure 1. This means that the landmark matrices $\tilde{Q}$ and $\tilde{K}$ are updated after receiving ${n}/{m}$ input tokens, following the segment-means process used in [19]. In practice, this will cause the oldest landmarks $\tilde{q}_{\text{old}}$ and $\tilde{k}_{\text{old}}$ (computed by using tokens that have already been shifted out of the $Q$ and $K$ matrices) to be shifted out of the landmark matrices $\tilde{Q}$ and $\tilde{K}$, respectively. The newest landmarks $\tilde{q}_{\text{new}}$ and $\tilde{k}_{\text{new}}$, computed as the mean of the most recent ${n}/{m}$ tokens in the $Q$ and $K$ matrices, respectively, will then be included in the landmark matrices $\tilde{Q}$ and $\tilde{K}$. As such, the update of the landmark matrices takes the form:
@@ IV-A1 Continual Inference with updated landmarks
where $B_{\phi}=(\phi(B)^{-1}\odot B)$ and $\Gamma_{\phi}=(\phi(\Gamma)^{-1}\odot\Gamma)$.
Following the Continual Transformer approach [30], a Single Output version of the Continual SDA is also proposed, as there are cases where only the attended features from the newest input token are needed, allowing for further reducing the number of computations. In practice, this is achieved by using the last row of the matrix $B_{\phi}$ rather than the full matrix i.e.:
@@
where $(\beta_{\phi})_{\text{new}}$ is the last row of the $B_{\phi}$ matrix. This still requires performing the common intermediate computations in the same way as its retroactive counterpart.
@@ IV-A2 Continual Inference with non-updated landmarks
where $\phi(\Delta)^{-1}_{\text{prev}}$ and $(\Delta V)_{\text{prev}}$ correspond to the matrices obtained in the previous inference step.
@@ IV-B Fixed Landmarks
After training the model and determining all landmarks, the SDA module used for Continual Retroactive Inference, i.e., $\text{Att}_{\text{CoNyRe}}^{\text{Fix}}$, has the form of Eq. (23), and the SDA module for the Continual Single Output Inference, i.e., $\text{Att}_{\text{CoNySi}}^{\text{Fix}}$, has the form of Eq. (26). The computational cost of this model is identical to the Nyström-based Continual Inference with non-updated landmarks for both the Retroactive and Single Output versions.
@@ IV-C Implementation Aspects
  • For training, we use a modified version of the non-continual model with the circular positional encoding described above and the corresponding landmark selection scheme as described in Sections IV-A and IV-B, depending on whether continual or fixed landmarks are used, respectively. We follow this approach as the non-continual training processes are faster when the entire sequence is available from the beginning, and both continual and non-continual SDA variants produce identical results.
@@
For replacing the softmax with exponential operations in both the Continual Transformers and our Continual Nyströmformers numerical stability issues need to be considered, as the calculation of the exponential can be prone to overflow and underflow. This is addressed in most softmax implementations (e.g., [39]) by employing the so-called stable softmax variant [40] $\text{s}(x)_{i}=\text{exp}(x_{i}-C)/\sum_{j=1}^{n}\text{exp}(x_{j}-C)$. By setting $C$ to the maximum value in $x$ possible overflow and underflow issues are addressed, as at least one value will be higher than zero after the calculation of the exponential operations. However, this approach cannot be applied in the continual versions of SDA, as the maximum values of $x$ in the attention window can change at every inference step. This would cause constant updates in most matrices, increasing the cost very significantly. In our experiments described in the next section, this issue has not been observed. In case such a stability issue is observed, using a dropout layer or other type of normalization layer before the SDA module can address it.
@@ V Experiments
In this section, we provide comparisons of models using the proposed SDA formulations with models using the original SDA formulation [1], the Nyström-based SDA formulation [19] and the Continual SDA formulation [28] in terms of their memory and computational costs, as well as their performance on Audio Classification and Online Action Detection tasks. A summary of the properties of these models is provided in Table II.
TABLE II: Properties of the SDA modules used in our experiments.
@@
$\text{Att}_{\text{CoNy}}^{\text{Cont}}$ (Ours) $\checkmark$ $\checkmark$ Continual
$\text{Att}_{\text{CoNy}}^{\text{Fix}}$ (Ours) $\checkmark$ $\checkmark$ Fixed
@@ V-B Memory overhead experiments
The rest of the SDA formulations exhibit a similar behavior, as all of them require storing intermediate matrices with size $n\times d$, including the full matrices $Q$, $K$ and $V$. The SDA formulations $\text{Att}_{\text{CoNy}}$ and $\text{Att}^{\text{Fix}}_{\text{CoNy}}$ have the highest memory cost in this group because they require storing the same set of matrices at peak memory time. The SDA formulations $\text{Att}_{\text{Co}}$ and $\text{Att}^{\text{Cont}}_{\text{CoNy}}$ have a very similar memory costs. This is due to that both SDA formulations need to store the same number of matrices of size $n\times d$ at peak memory time.
TABLE III: Audio classification on the GTZAN dataset.
@@
Att 1 $93.12\pm 5.44$ 5.57 $\times$1
2 $94.32\pm 4.31$ 11.13 $\times$1
$\text{Att}_{\text{Co}}$ 1 $93.16\pm 5.39$ 0.07 $\times$80.26
2 $94.52\pm 4.05$ 0.39 $\times$48.26
$\text{Att}_{\text{Ny}}$-4 1 93.54±$\pm 4.26$ 0.42 $\times$13.17
2 93.92±$\pm 4.43$ 0.85 $\times$13.17
$\text{Att}_{\text{Ny}}^{\text{Fix}}$-4 1 $93.24\pm 5.30$ 0.37 $\times$14.98
2 $93.76\pm 5.14$ 0.74 $\times$14.98
$\text{Att}_{\text{CoNy}}^{\text{Cont}}$-4 (Ours) 1 $92.92\pm 5.43$ 0.01 $\times$509.66
2 $93.54\pm 4.91$ 0.11 $\times$102.57
$\text{Att}_{\text{CoNy}}^{\text{Fix}}$-4 (Ours) 1 $90.76\pm 5.52$ 0.005 $\times$1028
2 $92.36\pm 5.52$ 0.10 $\times$108.92
TABLE IV: Online Action Detection on the THUMOS14 dataset using a Temporal Segment Network pre-trained on the Kinetics-400 dataset or ActivityNet dataset.
@@
Att 1 $64.32\pm 0.66$ $98.27\pm 0.05$ $56.00\pm 0.71$ $97.39\pm 0.04$ 8.46 $\times$1
2 $\textbf{64.66}\pm 0.30$ $98.28\pm 0.06$ $\textbf{56.95}\pm 0.80$ $97.36\pm 0.06$ 16.92 $\times$1
$\text{Att}_{\text{Co}}$ 1 $63.89\pm 0.14$ $98.33\pm 0.05$ $55.69\pm 0.41$ $\textbf{97.43}\pm 0.04$ 0.20 $\times$42.99
2 $63.93\pm 0.43$ $\textbf{98.34}\pm 0.02$ $56.04\pm 0.27$ $97.40\pm 0.08$ 0.65 $\times$25.88
$\text{Att}_{\text{Ny}}$-16 1 $59.74\pm 0.29$ $97.72\pm 0.03$ $50.57\pm 0.52$ $96.63\pm 0.05$ 4.71 $\times$1.80
2 $59.32\pm 0.23$ $97.72\pm 0.04$ $50.24\pm 0.14$ $96.47\pm 0.03$ 9.42 $\times$1.80
$\text{Att}_{\text{Ny}}^{\text{Fix}}$-16 1 $56.77\pm 0.51$ $97.46\pm 0.11$ $48.80\pm 0.75$ $96.41\pm 0.12$ 4.21 $\times$2.01
2 $53.80\pm 2.19$ $97.28\pm 0.29$ $48.53\pm 0.61$ $96.19\pm 0.11$ 8.43 $\times$2.01
$\text{Att}_{\text{CoNy}}^{\text{Cont}}$-16 (Ours) 1 $60.03\pm 0.33$ $97.70\pm 0.03$ $50.60\pm 0.13$ $96.64\pm 0.03$ 0.46 $\times$18.35
2 $59.30\pm 0.39$ $97.64\pm 0.07$ $50.68\pm 0.38$ $96.64\pm 0.04$ 1.43 $\times$11.84
$\text{Att}_{\text{CoNy}}^{\text{Fix}}$-16 (Ours) 1 $55.48\pm 0.26$ $97.21\pm 0.07$ $47.92\pm 0.69$ $96.05\pm 0.05$ 0.12 $\times$73.52
2 $52.41\pm 0.65$ $96.82\pm 0.13$ $46.60\pm 0.36$ $95.76\pm 0.06$ 1.26 $\times$13.40